find the area of the region bounded by the curve  y=ex & lines x=0 & y=e ?   

Grade:12

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Pratish,

 

The Req Area would be the area of the rectangle {from x=0 to x=1, with y=0 to y=e} - area bound by the curve y=e^x from x=0 to x=1.

So A = e - area of the curve

area of curve = Int(e^x) from 0 to 1 = e-1

So A = e - (e-1) = 1 sq unit.

 

Best Regards,

Ashwin (IIT Madras).